OpenStudy (anonymous):

What is the arc length if Θ = 6 pi over 5 and the radius is 2 cm?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

arc length is just \(r \theta = 2 \times 6\pi/5 = ...\)

OpenStudy (anonymous):

usE S=R*THETA here r=2cm and theta=6 pi /5 so S=(6 pi/5)*2=12pi/5=12*3.14/5=37.68/5=7.536cm

OpenStudy (anonymous):

ohhh k thanks :)